# On the enumeration of connected sets in finite lattice graphs

### v1

@@ Abstract
A connected set in a graph is a nonempty subset of vertices whose induced subgraph is connected. In an infinite lattice, a connected set is often referred to as a lattice animal, whose enumeration up to isomorphism is a classical problem in both combinatorics and statistical physics. In this paper, we focus on the enumeration of connected sets in finite lattice graphs, providing a link between combinatorial counting and structural connectivity in the system.
For any positive integers $m,n$, let $N(P_{m}\times P_{n})$ and $N(C_{m}\times P_{n})$ denote the number of all connected sets in the $(m\times n)$-lattice graph and $(m\times n)$-cylindrical lattice graph, respectively. In 2020, Vince derived enumeration formulas for $N(P_{m}\times P_{2})$ and $N(C_{m}\times P_{2})$, and highlighted the increasing difficulty of extending these calculation results to larger (cylindrical) lattice graphs. Recently, the authors of this paper have developed a method based on multi-step recurrence formulas to obtain enumeration results for $N(P_{m}\times P_{n})$ with $m\leq 4$. In this article, we apply a similar approach to derive the enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$. Further, for the general case, we establish an explicit and tight lower bound on the number of connected sets in any Cartesian product graph with the path of order $n$ as a factor, by employing the transfer matrix method on a subclass of connected sets. Based on this, corresponding asymptotic analysis is provided for some lattice graphs and other graphs of this type as $n$ tends to infinity.
@@ 1 Introduction
The enumeration of connected structures on lattices has long attracted interest in both combinatorics and statistical physics [10, 11, 15]. Examples include the dimer problem that counts perfect matchings [13, 23], and the counting of spanning forests and acyclic orientations [7, 1], independent sets [17, 2], and other graph parameters for geometric structures such as the Sierpiski gasket [4, 5, 6, 24]. In particular, lattice animals (finite sets of connected lattice sites or bonds) appear naturally in problems such as percolation, branching processes, and polymer networks as it models connected structures such as clusters and polymer configurations. Most research on the enumeration of lattice animals focuses on distinct animals of a given size on an infinite lattice up to translation and other symmetries, as well as the asymptotic behavior as the size of the animal tends to infinity.
@@
In this paper, we consider a related but complementary problem. Rather than focusing on an infinite lattice and distinct animals, we are interested in the number of all animals within a finite lattice region, which reflects the connectivity of the finite lattice regarding its size and geometry. This research is also motivated by, and originates from, relevant studies on the enumeration of spanning trees in lattices [8, 22]. From a statistical mechanics perspective, enumerating all animals in a finite region is similar to listing all microstates of a discrete system, with each animal corresponding to a distinct microconfiguration. In this way, our study connects combinatorial enumeration with finite-system statistical mechanics, linking structural counts directly to system connectivity and size effects. Furthermore, this approach may offer insights relevant to percolation and other lattice-based models in statistical physics.
@@
In general, for any graph $G=(V(G),E(G))$, a connected set $C$ of $G$ is a subset of $V(G)$ such that the subgraph of $G$ induced by $C$, denoted by $G[C]$, is connected. Clearly, when $G$ is a lattice, any connected set of $G$ is an animal and vice versa. Denote by $\mathcal{C}(G)$ the set of all the connected sets of $G$, let $N(G)=|\mathcal{C}(G)|$, and let $c(G)=\sqrt[|V(G)|]{N(G)}$. Note that $c(G)$ is a parameter that has been independently introduced or employed by many researchers to study the asymptotic behavior of $N(G)$ [3, 9, 12].
@@
In the literature, research on counting connected sets has been extensive and approached from various perspectives. For instance, Kangas et al. [12] determined upper bounds of $N(G)$ for graphs $G$ with the maximum dergee at most 5; Luo et al. [14] established sharp upper and lower bounds of $N(G)$ for any $G$ based on certain graph parameters; Cambie et al. [3] recently studied the maximum number of connected sets for regular graphs; Vince [19] conducted studies on several parameters derived from $N(G)$ and proved that $P_{n}$ minimizes the average order of connected sets over all connected graphs of order $n$; Haslegrave [9] bounded the connected set density away from 1 for graphs with the minimum degree at least 3; and Vince [20] showed that if $G$ is a connected vertex transitive graph, then the average order of the connected sets of $G$ is at least $|V(G)|/2$.
@@
In 2020, Vince [21] derived explicit enumeration formulas for connected sets in $P_{n}\times P_{2}$ and $C_{n}\times P_{2}$ that
$N(P_{n}\times P_{2})=\frac{\beta(n+3)-4n-7}{2}~~\text{and}~~N(C_{n}\times C_{2})=1-3n+2\beta(n)+3n\overline{\beta}(n),$
@@
As a preliminary study to Problem 1.1, the authors of this paper have developed a recursive method in [16] to determine the numbers of connected sets in $(3\times n)$- and $(4\times n)$-lattice graphs, accompanied by the explicit enumeration formula for $N(K_{m}\times P_{n})$ (refer to Theorem 2.1) by employing the transfer matrix method, where $K_{m}$ is the complete graph of order $m$.
@@
For any graph $G$, we call a maximal set in $G\times P_{n}$ consisting of all the vertices with the same second coordinate a column. Then any column induces a copy of $G$ in $G\times P_{n}$, and two columns $X,Y$ are said to be consecutive if there exist $(t,u)\in X$ and $(t,v)\in Y$ such that $uv\in E(P_{n})$. Let $N_{L}(G\times P_{n})$ be the number of connected sets $C$ of $G\times P_{n}$ that induce connected sets in $G\times P_{n}[X\cup Y]$ for every pair of consecutive columns $X,Y$ of $G\times P_{n}$ with $X\cap C,Y\cap C\neq\emptyset$. Then it is easy to see that $N_{L}(G\times P_{n})$ is a natural lower bound for $N(G\times P_{n})$. Moreover, $N_{L}(G\times P_{n})$ is a tight lower bound and its explicit enumeration formula is as follows.
@@
The proof of Theorem 1.2 will be given in Section 4. By Theorem 1.2, we further provide in Section 4 an asymptotic analysis on the behavior of $N(P_{3}\times P_{n})$, $N(C_{4}\times P_{n})$, $N(C_{5}\times P_{n})$ and $N(K_{1,3}\times P_{n})$ with respect to parameters $c(P_{3}\times P_{n})$, $c(C_{4}\times P_{n})$, $c(C_{5}\times P_{n})$ and $c(K_{1,3}\times P_{n})$ as $n$ tends to infinity, where $K_{1,3}$ is the star graph of order four. Note that a similar analysis can also be done for other Cartesian product graphs $G\times P_{n}$, although the calculations become more time-consuming as $|V(G)|$ increases.
@@ 2 The enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$
We first fix a labeling of $V(K_{m}\times P_{n})$ for all $m,n$, which naturally induces a labeling of $V(C_{m}\times P_{n})$ and $V(P_{m}\times P_{n})$, and will be used throughout Sections 2 and 3. Assume that $V(K_{m}\times P_{n})=\{v_{i,j}:i\in\{1,2,...,m\},j\in\{1,2,...,n\}\}$, where for any $i_{1},i_{2}\in\{1,2,...,m\}$ and $j_{1},j_{2}\in\{1,2,...,n\}$, $v_{i_{1},j_{1}}v_{i_{2},j_{1}}\in E(K_{m}\times P_{n})$, and $v_{i_{1},j_{1}}v_{i_{1},j_{2}}\in E(K_{m}\times P_{n})$ if and only if $|j_{1}-j_{2}|=1$. See an example in Figure 2. Then for each $j=1,2,\dots,n$, $I_{j}=\{v_{i,j}:i\in\{1,2,...,m\}\}$ is a column, and we call it the $j$-th column. For any $S_{j}\subseteq I_{j}$ with $j\geq 2$, let $S_{j-1}$ be the set of vertices in $I_{j-1}$ that are adjacent to the vertices in $S_{j}$. Then $|S_{j}|=|S_{j-1}|$ clearly holds.
@@
Let $\mathcal{C}_{m,n}$ denote the subset of connected sets of $K_{m}\times P_{n}$ that contain at least one vertex from each column. Then it is easy to see that $N(K_{m}\times P_{n})=\sum\limits_{k=1}^{n}(n-k+1)\cdot|\mathcal{C}_{m,k}|.$ For any $i=1,2,\dots,m$, let $f_{m,n}^{i}$ denote the number of sets in $\mathcal{C}_{m,n}$ that contain a fixed set of $i$ vertices in the $n$-th column. Then as illustrated in Theorem 1.2 of [16], $|\mathcal{C}_{m,n}|=\sum\limits_{i=1}^{m}\binom{m}{i}f_{m,n}^{i}$, and $f_{m,n}^{i}$ can be derived by the recurrence formula that
@@
By Theorem 2.1, we shall now focus on the set $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$. Let $\mathcal{C}_{m,n}^{\prime}$ be the subset of $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$, where each element of $\mathcal{C}_{m,n}^{\prime}$ contains at least one vertex from every column of $K_{m}\times P_{n}$. Clearly, $\mathcal{C}_{m,n}^{\prime}\subseteq\mathcal{C}_{m,n}$. For any nonempty set $X\subseteq V(K_{m}\times P_{n})$, let
@@
Remark. Note that for Lemma 2.3 (3), the only exceptional case is characterized by the following: each component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$, $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X])<k(C_{m}\times P_{n-2}[X]),$ and there exist four components $R_{p},R_{i},R_{q},R_{j}$ in $C_{m}\times P_{n-2}[X]$ with $p<i<q<j$ such that $R_{p},R_{q}$ are in the same component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ while $R_{i},R_{j}$ are in other components. An example is shown in Figure 3. For such cases, it is difficult to find a recurrence formula to calculate $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)$. However, for the case $m\leq 7$, we have $k(C_{m}\times P_{n-2}[X])\leq 3$ for any $X\subseteq I_{n-2}$, which implies that the situation described above does not arise.
@@ 3 Calculation results on $N(C_{4}\times P_{n})$ and $N(C_{5}\times P_{n})$
Similarly, we can establish the recurrence formulas for $N(C_{5}\times P_{n})$. For any positive integer $n$, assume that $v_{i+5,n}=v_{i,n}$ for all $i=1,2,3,4,5$, and let
@@ 4 Asymptotic analysis
Proof of Theorem 1.2. We begin by proving (1.1). For each $k=1,2,\dots,n$, assume that $V(P_{k})=\{v_{1},v_{2},\dots,v_{k}\}$, where $v_{i}v_{i+1}\in E(P_{k})$ for all $i\in\{1,2,\dots,k-1\}$. Then the $j$-th column $I_{j}$ of $G\times P_{k}$ is clearly $\{(x,v_{j}):x\in V(G)\}$ for all $j\in\{1,2,\dots,k\}$.
Let $\mathcal{C}_{L}(k)$ be the set of sets $C$ in $\mathcal{C}(G\times P_{k})$ such that $C\cap I_{j}\neq\emptyset$ for each $j\in\{1,2,...,k\}$ and $(C\cap I_{j})\cup(C\cap I_{j+1})\in\mathcal{C}(G\times P_{k})$ for each $j\in\{1,2,...,k-1\}$. Then it suffices to show that $\mathbf{1}^{T}\cdot A^{k-1}\cdot\mathbf{1}=|\mathcal{C}_{L}(k)|$ for all $k\geq 2$.
In fact, we shall show by induction that for each $(i,j)$-entry in $A^{k-1}$, it is the number of sets $C$ in $\mathcal{C}_{L}(k)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{k}):q\in\phi(j)\}$. Then (1.1) is proven by the definition of $\phi$.
When $k=2$, the matrix $A$ clearly satisfies our requirements. Suppose the statement is true for $k=t$, where $t\geq 2$. Let $l_{i,r}$ be any $(i,r)$-entry in $A^{t-1}$. Then $l_{i,r}$ is the number of sets in $\mathcal{C}_{L}(t)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t}):q\in\phi(r)\}$. Now we consider the case for $k=t+1$. Each $(i,j)$-entry in the matrix $A^{t}$ is $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$, where $a_{r,j}=1$ if and only if $\{(p,v_{t}):p\in\phi(r)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}\in\mathcal{C}(G\times P_{t+1}).$ Consequently, $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$ is the number of sets in $\mathcal{C}_{L}(t+1)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}$. Hence the matrix $A^{t}$ satisfies the requirements in the statement and (1.1) is proven.
Next, we show that (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. It is clear that the equality in (1.1) holds when $G$ is complete or $n\in\{1,2\}$. Now, assume that the equality in (1.1) holds for a connected graph $G$ that is not complete and $n\geq 3$. Then there exist nonadjacent vertices $u$ and $v$ in $G$ connected by a path $ux_{1}x_{2}\dots x_{t}v$ in $G$. Let
@@
Since $n\geq 3$, it can be easily verified that $C\in\mathcal{C}(G\times P_{n})$ and $C\cap I_{1},C\cap I_{2}\neq\emptyset$. However, $C\cap(I_{1}\cup I_{2})\notin\mathcal{C}(G\times P_{n})$. Hence, $N_{L}(G\times P_{n})<N(G\times P_{n})$, a contradiction. As a result, (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. $\Box$
@@
    Since $A_{1}$ is symmetric with nonnegative entries, the limit
    $\lim_{n\to\infty}\left(\mathbf{1}^{T}A_{1}^{n-1}\mathbf{1}\right)^{1/n}$
    exists by the Gelfand theorem, and is equal to the largest eigenvalue $\lambda_{A_{1}}$ of $A_{1}$. Using Mathematica, we find that $\lambda_{A_{1}}\approx 10.5318$. Hence, we obtain the following bound for $c(C_{4}\times P_{n})$:
    $c(C_{4}\times P_{n})>\lambda_{A_{1}}^{1/4}\approx 1.8014.$
@@
    $c(C_{5}\times P_{n})>\lambda_{A_{2}}^{\frac{1}{5}}\approx 1.7877.$
@@
    In fact, both bounds in (i) and (ii) provide a good estimate, as shown in Figure 4(a) and Table 1 that illustrate the asymptotic behavior of $c(C_{4}\times P_{n})$ and $c_{1}=(N_{L}(C_{4}\times P_{n}))^{1/4n}=\sqrt[4n]{13n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{1}^{k-1}\mathbf{1})\cdot(n-k+1)}$, and Figure 4(b) and Table 2 that show the asymptotic behavior of $c(C_{5}\times P_{n})$ and $c_{2}=(N_{L}(C_{5}\times P_{n}))^{1/5n}=\sqrt[5n]{21n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{2}^{k-1}\mathbf{1})\cdot(n-k+1)}$, when $1\leq n\leq 800$.
@@
    Then by using Mathematica, we have $\lambda_{A_{3}}\approx 4.6524$ and the following bound for $c(P_{3}\times P_{n})$:
    $c(P_{3}\times P_{n})>\lambda_{A_{3}}^{1/3}\approx 1.6694.$
@@
    $c(K_{1,3}\times P_{n})>\lambda_{A_{4}}^{1/4}\approx 1.7288.$

### v2

# On the enumeration of connected sets in finite cylindrical lattice graphs

@@ Abstract
A connected set in a graph is a non-empty set of vertices that induces a connected subgraph. In an infinite lattice, a connected set is often referred to as a lattice animal, whose enumeration up to isomorphism is a classical problem in both combinatorics and statistical physics. In this paper, we focus on the enumeration of connected sets in finite lattice graphs, providing a link between combinatorial counting and structural connectivity in the system.
For any positive integers $m,n$, let $N(P_{m}\times P_{n})$ and $N(C_{m}\times P_{n})$ denote the number of all connected sets in the $(m\times n)$-lattice graph $P_{m}\times P_{n}$ and $(m\times n)$-cylindrical lattice graph $C_{m}\times P_{n}$, respectively. In 2020, Vince derived enumeration formulas for $N(P_{m}\times P_{2})$ and $N(C_{m}\times P_{2})$, and highlighted the increasing difficulty of extending these calculation results to larger (cylindrical) lattice graphs. Recently, the authors of this paper have developed a method based on multi-step recurrence formulas to obtain the enumeration formula for $N(P_{m}\times P_{n})$ with $m\leq 4$. In this article, we apply a similar approach to derive the enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$. Further, for the general case, we establish an explicit and tight lower bound on the number of connected sets in the Cartesian product graph $G\times P_{n}$ for any connected graph $G$, by employing the transfer matrix method on a subclass of connected sets. Based on this, we perform an asymptotic analysis on several lattice graphs and show that $O(N(P_{3}\times P_{n}))=1.6694^{3n}$, $O(N(C_{4}\times P_{n}))=1.8014^{4n}$, and $O(N(C_{5}\times P_{n}))=1.7877^{5n}$.
@@ 1 Introduction
The enumeration of local structures in lattices has long been a topic of interest in both combinatorics and statistical physics. Classical examples include the dimer problem, which concerns the counting of perfect matchings [10, 20], as well as the enumeration of independent sets [14, 2], spanning forests and acyclic orientations [4, 1]. In particular, the enumeration of lattice animals (finite sets of connected lattice sites or bonds) arises naturally in studies of percolation, branching processes, and polymer networks, as these animals model connected structures such as clusters or polymer configurations. Most research in this area concentrates on counting distinct animals of a given size in an infinite lattice up to translation and other symmetries, and on analyzing the asymptotic behavior of these counts as the size of the animals tends to infinity [12, 8, 7]. While these studies provide valuable insights into large-scale combinatorial behavior, many mysteries about lattices remain.
@@
In this paper, we consider a related but complementary problem. Rather than focusing on an infinite lattice and distinct animals, we study the number of all animals within a finite lattice region, which reflects the connectivity of the finite lattice regarding its size and geometry. This research is also motivated by and originates from relevant studies on the enumeration of spanning trees in finite lattices [5, 19]. From a statistical mechanics perspective, enumerating all animals in a finite region is similar to listing all microstates of a discrete system, with each animal corresponding to a distinct microconfiguration. In this way, our study connects combinatorial enumeration with finite-system statistical mechanics, linking structural counts directly to system connectivity and size effects. Furthermore, this approach may offer insights relevant to percolation and other lattice-based models in statistical physics.
@@
For the remainder of this paper, we consider finite graphs only. For any graph $G=(V(G),E(G))$, a connected set $C$ of $G$ is a non-empty subset of $V(G)$ such that the subgraph of $G$ induced by $C$, denoted by $G[C]$, is connected. Clearly, when $G$ is a lattice, any connected set of $G$ is an animal and vice versa. Let $\mathcal{C}(G)$ be the set of all connected sets of $G$, let $N(G)=|\mathcal{C}(G)|$, and let $c(G)=\sqrt[|V(G)|]{N(G)}$. Then $N(G)$ is the number of all connected sets in $G$, and $c(G)$ is a parameter characterizing the asymptotic behavior of $N(G)$, where $1<c(G)<2$ as shown in [3, 6, 9].
@@
In the literature, research on counting connected sets has been extensive and approached from various perspectives. For instance, Kangas et al. [9] determined upper bounds of $N(G)$ for graphs $G$ with the maximum degree at most 5; Luo et al. [11] established sharp upper and lower bounds for $N(G)$ in terms of graph parameters such as the chromatic number, stability number, matching number and connectivity; Cambie et al. [3] recently studied the maximum number of connected sets among all $d$-regular graphs for small $d$; Vince [16] conducted studies on several parameters derived from $N(G)$ and proved that $P_{n}$ minimizes the average order of connected sets over all connected graphs of order $n$; Haslegrave [6] proved that the average connected set density is bounded away from 1 for graphs without vertices of degree 2; and Vince [17] showed that if $G$ is a connected vertex transitive graph, then the average order of the connected sets of $G$ is at least $|V(G)|/2$.
@@
In 2020, Vince [18] derived the following explicit enumeration formulas for connected sets in $P_{n}\times P_{2}$ and $C_{n}\times P_{2}:$
$N(P_{n}\times P_{2})=\frac{\beta(n+3)-4n-7}{2}~~\text{and}~~N(C_{n}\times P_{2})=1-3n+2\beta(n)+3n\overline{\beta}(n),$
@@
As a preliminary study to Problem 1.1, the authors of this paper have developed a recursive method in [13] to determine the numbers of connected sets in $(3\times n)$- and $(4\times n)$-lattice graphs, accompanied by the explicit enumeration formula for $N(K_{m}\times P_{n})$ (see Theorem 2.1) by employing the transfer matrix method, where $K_{m}$ is the complete graph of order $m$.
@@
For any connected graph $G$, we call a set in $G\times P_{n}$ consisting of all the vertices with the same second coordinate a column. Then any column induces a copy of $G$ in $G\times P_{n}$, and two columns $X,Y$ are said to be consecutive if there exist $(t,u)\in X$ and $(t,v)\in Y$ such that $uv\in E(P_{n})$. In the following, we shall focus on a special class of connected sets $\mathcal{C}_{L}(G\times P_{n})$, which is defined as $\{C\in\mathcal{C}(G\times P_{n}):C\cap(X\cup Y)\in\mathcal{C}(G\times P_{n})\text{ for any two consecutive columns }X,Y~\text{of}~G\times P_{n}~\text{such that}~X\cap C,Y\cap C\neq\emptyset\}$. Such sets are more densely connected, as illustrated in Figure 2.
[0.65]Refer to caption
(a) $C\in\mathcal{C}_{L}(C_{5}\times P_{4})$           (b) $C\notin\mathcal{C}_{L}(C_{5}\times P_{4})$
Figure 2: Examples for sets $C$ in $\mathcal{C}(C_{5}\times P_{4})$, where $C$ is the set of solid vertices
Now let $N_{L}(G\times P_{n})=|\mathcal{C}_{L}(G\times P_{n})|$. Then $N_{L}(G\times P_{n})\leq N(G\times P_{n})$ as $\mathcal{C}_{L}(G\times P_{n})\subseteq\mathcal{C}(G\times P_{n})$. Moreover, $N_{L}(G\times P_{n})$ is a tight lower bound of $N(G\times P_{n})$ and its explicit enumeration formula is as follows.
@@
The proof of Theorem 1.2 will be given in Section 4.
By Theorem 1.2, we further provide in Section 4 an asymptotic analysis on the behavior of $N(P_{3}\times P_{n})$, $N(C_{4}\times P_{n})$, $N(C_{5}\times P_{n})$ and $N(K_{1,3}\times P_{n})$ as $n$ tends to infinity, where $K_{1,3}$ is the star graph of order four. Note that a similar analysis can also be done for other Cartesian product graphs $G\times P_{n}$, although the calculations become more time-consuming as $|V(G)|$ increases.
@@ 2 The enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$
We first fix a labeling of $V(K_{m}\times P_{n})$ for all $m,n$, which naturally induces a labeling of $V(C_{m}\times P_{n})$ and will be used throughout Sections 2 and 3. Assume that $V(K_{m}\times P_{n})=\{v_{i,j}:i\in\{1,2,...,m\},j\in\{1,2,...,n\}\}$, where for any $i_{1},i_{2}\in\{1,2,...,m\}$ and $j_{1},j_{2}\in\{1,2,...,n\}$, $v_{i_{1},j_{1}}v_{i_{2},j_{1}}\in E(K_{m}\times P_{n})$, and $v_{i_{1},j_{1}}v_{i_{1},j_{2}}\in E(K_{m}\times P_{n})$ if and only if $|j_{1}-j_{2}|=1$. See an example in Figure 3. Then for each $j=1,2,\dots,n$, $I_{j}=\{v_{i,j}:i\in\{1,2,...,m\}\}$ is a column, and we call it the $j$-th column. For any $S_{j}\subseteq I_{j}$ with $j\geq 2$, let $S_{j-1}$ be the set of vertices in $I_{j-1}$ that are adjacent to the vertices in $S_{j}$. Then $|S_{j}|=|S_{j-1}|$ clearly holds.
@@
Let $\mathcal{C}_{m,n}$ denote the subset of $\mathcal{C}(K_{m}\times P_{n})$ that contain at least one vertex from each column. Then it is easy to see that $N(K_{m}\times P_{n})=\sum\limits_{k=1}^{n}(n-k+1)\cdot|\mathcal{C}_{m,k}|.$ For any $i=1,2,\dots,m$, let $f_{m,n}^{i}$ denote the number of sets in $\mathcal{C}_{m,n}$ that contain a fixed set of $i$ vertices in the $n$-th column. Then as illustrated in Theorem 1.2 of [13], $|\mathcal{C}_{m,n}|=\sum\limits_{i=1}^{m}\binom{m}{i}f_{m,n}^{i}$, and $f_{m,n}^{i}$ can be derived by the recurrence formula that
@@
By Theorem 2.1, we shall now focus on the set $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$. Let $\mathcal{C}_{m,n}^{\prime}$ be the subset of $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$, where each element of $\mathcal{C}_{m,n}^{\prime}$ contains at least one vertex from every column of $K_{m}\times P_{n}$. Clearly, $\mathcal{C}_{m,n}^{\prime}\subseteq\mathcal{C}_{m,n}$. For any non-empty set $X\subseteq V(K_{m}\times P_{n})$, let
@@
Remark. Note that for Lemma 2.3 (3), the only exceptional case is characterized by the following: each component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$, $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X])<k(C_{m}\times P_{n-2}[X]),$ and there exist four components $R_{p},R_{i},R_{q},R_{j}$ in $C_{m}\times P_{n-2}[X]$ with $p<i<q<j$ such that $R_{p},R_{q}$ are in the same component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ while $R_{i},R_{j}$ are in other components. An example is shown in Figure 4. For such cases, it is difficult to find a recurrence formula to calculate $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)$. However, for the case $m\leq 7$, we have $k(C_{m}\times P_{n-2}[X])\leq 3$ for all $X\subseteq I_{n-2}$, which implies that the situation described above does not arise.
@@ 3 Calculation results on $N(C_{4}\times P_{n})$ and $N(C_{5}\times P_{n})$
Similarly, we establish the recurrence formulas for $N(C_{5}\times P_{n})$. For any positive integer $n$, assume that $v_{i+5,n}=v_{i,n}$ for all $i=1,2,3,4,5$, and let
@@ 4 Asymptotic analysis
Proof of Theorem 1.2. We begin by proving (1.1). For each $k=1,2,\dots,n$, assume that $V(P_{k})=\{v_{1},v_{2},\dots,v_{k}\}$, where $v_{i}v_{i+1}\in E(P_{k})$ for all $i\in\{1,2,\dots,k-1\}$. Then the $j$-th column $I_{j}$ of $G\times P_{k}$ is clearly $\{(x,v_{j}):x\in V(G)\}$ for all $j\in\{1,2,\dots,k\}$. Moreover, let
$\mathcal{S}_{L}(k)=\{C\in\mathcal{C}_{L}(G\times P_{k}):C\cap I_{j}\neq\emptyset~\text{for each}~j=1,2,...,k\}.$
Then it suffices to show that $\mathbf{1}^{T}\cdot A^{k-1}\cdot\mathbf{1}=|\mathcal{S}_{L}(k)|$ for all $k\geq 2$.
In fact, we shall show by induction that for each $(i,j)$-entry in $A^{k-1}$, it is the number of sets in $\mathcal{S}_{L}(k)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{k}):q\in\phi(j)\}$. Then (1.1) is proven by the definition of $\phi$.
When $k=2$, the matrix $A$ clearly satisfies our requirements. Suppose the statement is true for $k=t$, where $t\geq 2$. Let $l_{i,r}$ be any $(i,r)$-entry in $A^{t-1}$. Then $l_{i,r}$ is the number of sets in $\mathcal{S}_{L}(t)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t}):q\in\phi(r)\}$. Now we consider the case for $k=t+1$. Each $(i,j)$-entry in the matrix $A^{t}$ is $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$, where $a_{r,j}=1$ if and only if $\{(p,v_{t}):p\in\phi(r)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}\in\mathcal{C}(G\times P_{t+1}).$ Consequently, $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$ is the number of sets in $\mathcal{S}_{L}(t+1)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}$. Hence the matrix $A^{t}$ satisfies the requirements in the statement and (1.1) is proven.
Next, we show that (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. It is clear that the equality in (1.1) holds when $G$ is complete or $n\in\{1,2\}$. Now, assume that the equality in (1.1) holds for a connected graph $G$ that is not complete and $n\geq 3$. Then there exist non-adjacent vertices $u$ and $v$ in $G$ connected by a path $ux_{1}x_{2}\dots x_{t}v$ in $G$. Let
@@
Since $n\geq 3$, it can be easily verified that $C\in\mathcal{C}(G\times P_{n})$ and $C\cap I_{1},C\cap I_{2}\neq\emptyset$. However, $C\cap(I_{1}\cup I_{2})\notin\mathcal{C}(G\times P_{n})$, implying that $C\in\mathcal{C}(G\times P_{n})\setminus\mathcal{C}_{L}(G\times P_{n})$. Hence $N_{L}(G\times P_{n})<N(G\times P_{n})$, a contradiction. As a result, (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. $\Box$
@@
    By the Gelfand theorem and the fact that $A_{1}$ is symmetric with all non-negative entries, we have
    $\lim_{n\to\infty}\left(\mathbf{1}^{T}A_{1}^{n-1}\mathbf{1}\right)^{1/n}=\lambda_{A_{1}},$
    where $\lambda_{A_{1}}$ is the largest eigenvalue of $A_{1}$. Using Mathematica, we find that $\lambda_{A_{1}}\approx 10.5318$. Hence for sufficiently large $n$,
    $c(C_{4}\times P_{n})=N(C_{4}\times P_{n})^{1/4n}>\left(\mathbf{1}^{T}A_{1}^{n-1}\mathbf{1}\right)^{1/4n}=\lambda_{A_{1}}^{1/4}\approx 1.8014.$
@@
    $c(C_{5}\times P_{n})=N(C_{5}\times P_{n})^{1/5n}>\lambda_{A_{2}}^{1/5}\approx 1.7877.$
@@
    Moreover, $c_{1}=(N_{L}(C_{4}\times P_{n}))^{1/4n}$ and $c_{2}=(N_{L}(C_{5}\times P_{n}))^{1/5n}$ provide a good estimate of $c(C_{4}\times P_{n})$ and $c(C_{5}\times P_{n})$, respectively. See some data in Figure 5, Table 1 and Table 2 as examples that illustrate the asymptotic behavior of $c(C_{4}\times P_{n})$, $c(C_{5}\times P_{n})$, $c_{1}$ and $c_{2}$.
  4. (iv)
    Additionally, based on extensive Mathematica computations, we suspect that the limits of $c(C_{4}\times P_{n})$ and $c(C_{5}\times P_{n})$ exist as $n$ tends to infinity, with approximate values 1.8374 and 1.8670, respectively.
@@
    By using Mathematica, $\lambda_{A_{3}}\approx 4.6524$, and
    $c(P_{3}\times P_{n})=N(P_{3}\times P_{n})^{1/3n}>\lambda_{A_{3}}^{1/3}\approx 1.6694.$
@@
    $c(K_{1,3}\times P_{n})=N(K_{1,3}\times P_{n})^{1/4n}>\lambda_{A_{4}}^{1/4}\approx 1.7288.$